A right triangular plate of base 2.0 m and height 1.0 m is submerged vertically in water, with top vertex 3.0 m below the surface. Find the force on one side of the plate.
1
Expert's answer
2020-04-13T19:07:39-0400

Given,

Base ="2.0 \space m"

Height = "1.0 \space m"

We know Density of water = "w=9800 N\/m^3"

Force = "?"

Here, x = length =

y = depth = "1"

So, ordered pair ="(0,1)"

One more ordered pair = "(2, 4)"

Equation of a line is "y = mx+ c"

Plug "(0, 1)"

"1 = 0 +c \\
c= 1 \\"

The equation becomes "y=mx+ 1"


Plug "(2, 4)" in the above equation


"4 = 2m + 1 \\
2m = 3 \\
m = \frac {3}{2}"

so the equation is "y =\frac {3}{2} x + 1"


So, "x = \frac {2y-2}{3}"


Force ="w \int_a ^b x \space y \space dy"



"= 9800 \int_1 ^4 \frac {2y-2}{3} y \space dy"



"= \frac {9800 \times 2}{3} \int _1 ^4 ( y^2 - y) dy"

"= \frac {9800 \times 2}{3} ( \frac {y^3} {3} - \frac {y^2}{2})_1 ^4"

"= \frac {19600} {3} (\frac {64}{3} - \frac {16} {2} - \frac {1}{3} + \frac {1}{2} )"


"= \frac {19600} {3} (\frac {128}{6} - \frac {48} {6} - \frac {2}{6} + \frac {3}{6} )"

"= \frac {19600} {3} (\frac {128-48-2+3}{6} ) =\frac {19600} {3} (\frac {81}{6})"

"Force = 88200 \space N"